# Approximate $SU(5)$, Fine Structure Constants

### v1

@@ Abstract
We fit the three finestructure constants of the Standard Model with three, in first approximation theoretically estimable parameters, 1) a “unified scale”,turning out not equal to the Planck scale and thus only estimable by a very speculative story, 2) a “number of layers” being a priori the number of families, and 3) a unified coupling related to a critical coupling on a lattice. So formally we postdict the three fine structure constants!
In the philosophy of our model there is a physically lattice theory with link variables taking values in a (or in the various) “small” representations of the Standard Model Group. We argue for that these representations function in first approximation as were the theory a genuine $SU(5)$ theory. Next we take into account fluctuation of the gauge fields in the lattice and obtain a correction to the a priori $SU(5)$ approximation, because of course the link fluctuations not corresponding any Standard model Lie algebra, but only to the SU(5), do not exist.
The model is a development of our old anti-grand-unification model having as its genuine gauge group, close to fundamental scale, a cross product of the standard model group $S(U(3)\times U(2))$ with itself, there being one Cartesian product factor for each family.
In these old works we included the hypotesis of “multiple point criticallity principle” which here effectively means the coupling constants be critical on the lattice. Counted relative to the Higgs scale we suggest the in our sense “unified scale” (where the deviations between the inverse fine structure constants deviate by quantum fluctuations being only from standard model groups, not SU(5)) makes up the 2/3 th power of the Planck scale relative to the Higgs scale, or better the topquarkmass scale..
@@ 1 Introduction
but mainly the Abelian coupling of $U(1)$ was not so well predicted contrary to the non-abelian ones (the attempt by Don Bennet and myself [14] got good numbers, but the theory is a bit complicated). Further Laperashvili and Das and Ryzhikh [16, 21, 5] have even united this type of model with grand unification with $SU(5)$ [16, 21]. They used also supersymmetry in their picture.
@@ 1.1 Character of Our Prediction(s)
The main point of the present article is really to predict the deviation from exact $SU(5)$ GUT at a certain scale $\mu_{U}$ at which we calculate the corrections to the exact $SU(5)$ inverse fine structure constants in the standard model as due to quantum fluctuations in the lattice theory assumed to be really physically existent at some scale. Since we predict the absolute values of the differences between the inverse finestructure constants at the scale, we have at this scale two numerical predictions, and thus we can afford to use one of these predictions at the fundamental scale to fit the scale, and we shall still have one predition left. For instance we can use the prediction at the scale, at which the ratio of the difference $1/\alpha_{2}(\mu_{U})-1/\alpha_{1\;SU(5)}(\mu_{U})$ to the other difference $1/\alpha_{1\;SU(5)}(\mu_{U})-1/\alpha_{3}(\mu_{U})$ shall be 2 to 3 (as our calculation implies). This is illustrated on figure 1, and one shall remark, that the three crossings of the inverse fine structure constant with the vertical black line on the figure at the scale about $5.*10^{13}GeV$ has been fitted, so that the three crossings lie with the ratio 2 : 3 of the two intervals. The $U(1)$ inverse fine structure constant passes in between the $SU(2)$, above it with a piece that is proportional to 2, and the $SU(3)$ line, then below it with a distance proportional to 3. But having fixed the scale $\mu_{U}$ this way it is still a very nontrivial prediction that e.g. the absolute difference between the $SU(2)$-crossing and the $SU(3)$-crossing is just $4.7123853\pi/2=$ 4.712385. This is illustrated on figure 2.
@@ 1.3 Plan of Article
Finally in section 12 we conclude, but also include some thoughts about the problems or suggestion for a quantum gravity, if we take the present work so serious, that we must claim that the fundamental scale for the Standard Model is the unification one, for our approximate $SU(5)$ GUT, even a bit low in energy compared to the usual unified scale. A lattice, which fluctuates even in scale in some background of a manifold or a projective space. If one could have the lattice imbedded in the continuum space wtih some symmetry including scalings, there might be a chanse of having a different way to average over fluctuations in the lattice constant size (i. e. coping with a fluctuating “fundamental scale”) for the fine structure constants gauge theories and for the gravity. Such different averaging can seperate the different scales to be observed for the two groups of forces,the Standard Model ones , and gravity.
@@ 4 Wilson Action
We here used that the plaquette variables, say $H^{a}(\Box)$ for the different plaquettes $\Box$ are not independently integrated over. On the contrary for each cube in the lattice there is a constraint which linearized means that the sum of six plaquaette variable for the plaquettes around the cube is restricted to be zero. Since there in 4 dimensions are 6 plaquettes per site and 4 cubes, this restriction would in first go mean that there per site were only 2 independent plaquette variables, but that is, however, not true, because there is a constraint between the four cube-constraints on the plaquettes. So in reality there is per site 3 independent constrains on 6 a priori plaquette variables. This gives that the average of the square $(H^{a})^{2}$ of a (Gaussian distributed) plaquaette variable get reduced by a factor 6 to (6-3) meaning a factor 2. Simplifying to just 2 variables to get restricted to 1 independent we could just think of a Gaussian distribution about the origo in a plane, and that we then restrict the at first two dimensional to a diagonal - a single dimension - being a restriction symmetric between the two orinal variables thought of as the coordinates. Then the restricted distribtuion on the symmetric diagonal would project into one of the coordinate axes with the average of the saquare diminished by a factor 2.
@@ 4.2 On Table
Use the numbers from thee table being just traces of the products of the diagonal matrices, which are noramlized so that their traces are $1$ for the $h^{2}$ and the dimension of the Lie Group for the $H_{i}^{2}$ normalizes the difference $\frac{1}{\alpha_{2}}-\frac{1}{\alpha_{3}}$ to one “unit” ignoring yet the factor 3 of number of families, and the hereby absorbed denominator 2, being
@@
now in the notation with “Re Tr” (in which it would at first have been $\pi$. So the prediction will be that the difference at the unifying scale of the two nonabelian inverse fine strucutre constants - which had number 1 (when the explicit 1/2 not included - will be $\frac{\pi}{2}$ for only one family, but $23*\pi/$ 2 for three families.
@@ 5.1 Tadpole correction calculations
But there is a further point in extracting the fine structure constant used in the work by Nyaizi et al: They use L’́uscher-Weisz action which even in the large $\beta=\beta_{pl}$ limit has an extra term consisting of double plachette actions with a coefficient which according to [33] is given by the
@@ 6 Fitting
The first step in our fitting of our model is to calculate the “unifying” scale $\mu_{u}$, at which the ratios between the differences between the inverse fine structure constants for the three subgroups of the Standard Model group is the one predicted from our calculation of the quantum fluctuation corrections. In fact the three inverse fine structure constants shall lie on the number axis as the numbers $(2,13/5,3)$ corresponding to the subgroups $(SU(3),U(1),SU(2)$, where we have chosen the $SU(5)$-normalisation for the $U(1)$-finstructure constant. The relation is expressed in terms of the two independent differences, that can be formed. Let us , e.g., say
@@
$29.57\pm 0.06\%-\frac{5}{3}*59.00\pm 0.02\%+\frac{2}{3}*8.446\pm 0.6\%$ $=$ $*\displaystyle\frac{-19/6-5/3*41/10+2/3*(-7)}{2\pi}*$
** $\ln\frac{\mu_{u}}{M_{Z}}$
@@ 7 What says our result about Original variables?
$*\displaystyle-\frac{b_{2}-b_{3}}{b_{2}-5/3*b_{1}+2/3b_{3}}*$
*$*(1/\alpha_{2}(M_{Z})-5/3*1/\alpha_{1\;SU(5)}(M_{Z})+2/3*1/\alpha_{3}(M_{Z})).$
@@ 8.2 Superfluous case Difference 1 to 3
Although it is just related to the twoo foregoing let us explicitely for check also calculate what the our requirement for the difference 1to 3 means
@@ 9.1 More thoughts on the Critical Coupling, Unified Coupling
Thinking a bit deeper: We should really not take a formula for the $SU(5)$ critical coupling without correction, because we have been claiming all through the article that in our model the $SU(5)$ symmetry and all its degrees and freedom do not exist. Rather we should look for correcting the number for the critical $\alpha_{5}$ to the critial coupling for the lattice standard Model group coupling:
@@ 10 Crude Second Order Calculation
One unneccesary ignorance of second order terms, which are expected to be of the order $(1/10)^{2}$ times the main term is, that we above let the $\alpha$ appearing as a factor in the $<H^{2}>$’s cancel with the $1/\alpha$ whichever among the $1/\alpha_{i}$’s we meet. Actually it was tempting to think, that by using this lucky trick of getting rid of the parameters in the estimate of our corrections, we were likely actually to get a better result with respect to agreeing for our calculations. Once we look for accuracies of the second order, there may be more corrections, such as the $H$ distribtuion being not even just Gaussian, and the whole program of doing the second order deserves a further article. Here we shall only make a very crude attempt to estimate the effect of seeing what $\alpha$ (among the three one say) comes into which of the fluctuations $<H^{2}>$. We shall make the assumption that the $\alpha$ to be used for an $H_{i}$ where it is the fuctuation in one of the basis vectors for the subgroup $i$ of the $SU(5)$, is $\alpha_{i}$. Then we see from table 1 that we have had relativley good luck by letting the two $\alpha$’s cancel each other, because the mostly contrbuting $<H_{i}^{2}>$ to the correction for the inverse fine structure constant $1/\alpha_{j}$, for the standard model subgroup denoted $j$ is actually mostly $i=j$ itself. In fact e.g. according to the table 1 the correction to
@@
For the $U(1)$ inverse fine structure constant the dominant cotrbutuion to the corrections comes from the two nonabelian groups, i.e. from $H_{2}$ and $H_{3}$, but it has a bigger number from the $H_{1}$ than any of the other two groups, namely $7/30$. But since the $U(1)$ coupling correction is so mixed, to take all the same $\alpha$ is not so bad.
@@ 11 Speculative Relation to Planck Scale
A major problem and surprize comming, if one takes our suggestion of truly existing lattice at the approximate or ours unification scale $\mu_{U}$ = $5.18*10^{13}GeV$ seriously is, that it suggests a “fundamental” scale quite different from the Planck scale. To seek a way out of this problem we propose to think of a fluctuating lattice even in size of the lattice constant in the sense that we speculate, that the general theory of relativity is still perturbatively treatable and rather well understood already - so that no completely speculated quatum gravity theory is needed at the $\mu_{U}$ scale - so that the whole lattice structure must be in a quantum superpostion state invariant under the reparametrization group from the general relativity. That is to say, with the philosophy, that there is very big quantum fluctuations in the gauge and taking the diffeomorphism of reparametrization symmetry as the gauge symmetry of general relativity, we must take it that the world is in a superposition of all the possible deformations of the lattice - needed for our model for the approximate GUT $SU(5)$ achieved by reparametrizations. That is to say, that in a typical component in this superpostion somewhere we find a very small lattice constant and somewherewe find a very big one, so that lattice cannot be exactly a Wilson one e.g.. But locally it could still be close to a Wilson lattice. Then of course the lattice constant value suggested by our parameter $\mu_{U}$ as lattice constant $a\approx 1/\mu_{U}$ could only be true in an average sense:
@@ 11.3 Ambiguity of Concept of Planck Energy Scale, reduced ?
$\hbox{and has }\ln(\frac{M_{Z}}{m_{av.ch.fermions}}$ $=$ 6.3276.327 6.3276.327 (273)
@@
In fact we think, we can argue for, that this latter choice is not the correct one, because the $8\pi$ or $4\pi$ usually comes from the differnce in the coefficient to a Coulomb field and the charge appearing in the field theory action. When we have just used the fermion masses without any $4\pi$-like correction we associate it with the simple relation $m=g_{y}\phi>$, while if I would like the Yukawa-field aruond the Higgs particle I would get a $1/(4\pi)$ factor in. So the simple masses correspond we could say to the Yukawa coupling $g_{y}$ being used for unit, and not the alternative $g_{y}/(4\pi)$. So to speak

### v3

@@ Abstract
We fit the three finestructure constants of the Standard Model with three, in first approximation theoretically estimable parameters, 1) a “unified scale”,turning out not equal to the Planck scale and thus only estimable by a very speculative story, 2) a “number of layers” being a priori the number of families, and 3) a unified coupling related to a critical coupling on a lattice. So formally we postdict the three fine structure constants! In the philosophy of our model there is a physically lattice theory with link variables taking values in a (or in the various) “small” representations of the Standard Model Group. We argue for that these representations function in first approximation as were the theory a genuine $SU(5)$ theory. Next we take into account fluctuation of the gauge fields in the lattice and obtain a correction to the a priori $SU(5)$ approximation, because of course the link fluctuations not corresponding any Standard model Lie algebra, but only to the SU(5), do not exist. The model is a development of our old anti-grand-unification model having as its genuine gauge group, close to fundamental scale, a cross product of the standard model group $S(U(3)\times U(2))$ with itself, there being one Cartesian product factor for each family. In these old works we included the hypotesis of “multiple point criticallity principle” which here effectively means the coupling constants be critical on the lattice. Counted relative to the Higgs scale we suggest the in our sense “unified scale” (where the deviations between the inverse fine structure constants deviate by quantum fluctuations being only from standard model groups, not SU(5)) makes up the 2/3 th power of the Planck scale relative to the Higgs scale, or better the topquarkmass scale..
Keywords
grand unfication SU(5); lattice theory; running couplings; anti-GUT; Standard Model Group;critical coupling; energy scales; fine structure constants; Planck scale
@@ 1 Introduction
but mainly the Abelian coupling of $U(1)$ was not so well predicted contrary to the non-abelian ones (the attempt by Don Bennett and myself [14] got good numbers, but the theory is a bit complicated). Further Laperashvili and Das and Ryzhikh [18, 23, 5] have even united this type of model with grand unification with $SU(5)$ [18, 23]. They used also supersymmetry in their picture.
@@
1.1 Some Previous attempts
Since it has long been wellknown that using minal SU(5) there is no working unifying scale in the sense that we see on e.g. fig 1 that the three lines representing the runnings of the three Standard Model Lie groups do not cross just in a point as SU(5) GUT would say. The two very popular ways of seeking to solve this problem are :
  • Super symmetry By introducing supersymmetry one has to have a scale for breaking this susy and thus of course a new parameter with which to adjust the crossing of the three lines. It happens and this is then a success for susy that this susy-breaking scalebeing very close where we have the data and thus to the place from which susy is possible. We namely expect from the hierarchy problem which susy can rescue that the breaking scale shall be very in energy. So susy like our model has a theoretical story about the value of parameter introduced to fit the SU(5). We have instead that our quantum correction shall be just the number of families times bigger than calculated with a simple lattice.
  • Bigger Groups containing SU(5)
    If you have a group containing SU(5) as a subgroup such as SO(10) there are possibilities for having the three standard model groups so to speak be packed into bigger simple groups at different ranges of scales and thus also get a way to achieve an extra parameter and thus solve the problem.
    In our philosophy we start from the group of the Standard Model which is $S(U(2)\times U(3))$, which means we already have the favoured group fixed, and then argue that the “smallest” / essentially simplest representation to use to construct the plaquette variable action contribtuion happens to be an $SU(5)$ representation. That is to say we claim that taking the Standard Model very seriously there is pointing to the group $SU(5)$ proper, in a way we do not have for the alternative unification groups.
    This is of course a very weak argument in favour of just SU(5) only, and thus in theories, that have some prediction for the energy scales of the various symmetry breakings, a prediction of the minimal SU(5)-breaking in the fine structure constants would be of an anlogous strength as our somewhat ad hoc factor 3 on the quantum correction fromthe lattice.
In general one should bring it as a propaganda for our model, that apart from the lattice - which should at the end be assumed to be fluctuating in link size from place to place in Minkowski space as quantum fluctuations -, a lattice that is somehow needed to avoid divergencies, we keep to just the Standard Model and nothing else, at least if you ignore our story about the critical coupling, which is only important for settling the unified fine structure constant. In other words our model claims to be more minimal than the just mentioned competing scenarious susy and grand unification with larger than SU(5) groups,or even with true SU(5) full symmetry.
@@ 1.2 Character of Our Prediction(s)
The main point of the present article (recently further developped in [16, 15]) is really to predict the deviation from exact $SU(5)$ GUT at a certain scale $\mu_{U}$ at which we calculate the corrections to the exact $SU(5)$ inverse fine structure constants in the standard model as due to quantum fluctuations in the lattice theory assumed to be really physically existent at some scale. Since we predict the absolute values of the differences between the inverse finestructure constants at the scale, we have at this scale two numerical predictions, and thus we can afford to use one of these predictions at the fundamental scale to fit the scale, and we shall still have one predition left. For instance we can use the prediction at the scale, at which the ratio of the difference $1/\alpha_{2}(\mu_{U})-1/\alpha_{1\;SU(5)}(\mu_{U})$ to the other difference $1/\alpha_{1\;SU(5)}(\mu_{U})-1/\alpha_{3}(\mu_{U})$ shall be 2 to 3 (as our calculation implies). This is illustrated on figure 1, and one shall remark, that the three crossings of the inverse fine structure constant with the vertical black line on the figure at the scale about $5.*10^{13}GeV$ has been fitted, so that the three crossings lie with the ratio 2 : 3 of the two intervals. The $U(1)$ inverse fine structure constant passes in between the $SU(2)$, above it with a piece that is proportional to 2, and the $SU(3)$ line, then below it with a distance proportional to 3. But having fixed the scale $\mu_{U}$ this way it is still a very nontrivial prediction that e.g. the absolute difference between the $SU(2)$-crossing and the $SU(3)$-crossing is just $4.7123853\pi/2=$ 4.712385. This is illustrated on figure 2.
@@ 1.4 Plan of Article
Finally in section 12 we conclude, but also include some thoughts about the problems or suggestion for a quantum gravity, if we take the present work so serious, that we must claim that the fundamental scale for the Standard Model is the unification one, for our approximate $SU(5)$ GUT, even a bit low in energy compared to the usual unified scale. A lattice, which fluctuates even in scale in some background of a manifold or a projective space. If one could have the lattice imbedded in the continuum space with some symmetry including scalings, there might be a chanse of having a different way to average over fluctuations in the lattice constant size (i. e. coping with a fluctuating “fundamental scale”) for the fine structure constants gauge theories and for the gravity. Such different averaging can seperate the different scales to be observed for the two groups of forces,the Standard Model ones , and gravity.
@@ 2 Our Model
The symbol ${\bf Z}_{app}$ denotes a discrete subgroup of the center of the covering group ${\bf R}\times SU(2)\times SU(3)$ which is determined by the Lie algebra of the Standard Model. The crucial requirement is that the elements is this subgroup of the center shall be represented trivially in all the representation for the various fields/particels in the Standard Model. This requirement leads to a rule ensuring in fact that leptons must have integer electric charges and and that the quarks must have charges 1/3 modulo 1. This means that some of the victories usually ascribed to SU(5) grand unification here is considered built into the global group.
@@ 4 Wilson Action
We here used that the plaquette variables, say $H^{a}(\Box)$ for the different plaquettes $\Box$ are not independently integrated over. On the contrary for each cube in the lattice there is a constraint which linearized means that the sum of six plaquaette variable for the plaquettes around the cube is restricted to be zero. Since there in 4 dimensions are 6 plaquettes per site and 4 cubes, this restriction would in first go mean that there per site were only 2 independent plaquette variables, but that is, however, not true, because there is a constraint between the four cube-constraints on the plaquettes. So in reality there is per site 3 independent constrains on 6 a priori plaquette variables. This gives that the average of the square $(H^{a})^{2}$ of a (Gaussian distributed) plaquaette variable get reduced by a factor 6 to (6-3) meaning a factor 2. Simplifying to just 2 variables to get restricted to 1 independent we could just think of a Gaussian distribution about the origo in a plane, and that we then restrict the at first two dimensions to a diagonal - a single dimension - being a restriction symmetric between the two orinal variables thought of as the coordinates. Then the restricted distribtuion on the symmetric diagonal would project into one of the coordinate axes with the average of the saquare diminished by a factor 2.
@@ 4.2 On Table
Use the numbers from thee table being just traces of the products of the diagonal matrices, which are normalized so that their traces are $1$ for the $h^{2}$ and the dimension of the Lie Group for the $H_{i}^{2}$ normalizes the difference $\frac{1}{\alpha_{2}}-\frac{1}{\alpha_{3}}$ to one “unit” ignoring yet the factor 3 of number of families, and the hereby absorbed denominator 2, being
@@
now in the notation with “Re Tr” (in which it would at first have been $\pi$). So the prediction will be that the difference at the unifying scale of the two nonabelian inverse fine strucutre constants - which had number 1 (when the explicit 1/2 not included) - will be $\frac{\pi}{2}$ for only one family, but $23*\pi/$ 2 for three families.
@@ 5.1 Tadpole Correction Calculations
But there is a further point in extracting the fine structure constant used in the work by Nyaizi et al: They use L’́uscher-Weisz action which even in the large $\beta=\beta_{pl}$ limit has an extra term consisting of double plaquette actions with a coefficient which according to [35] is given by the
@@ 6 Fitting
The first step in our fitting of our model is to calculate the “unifying” scale $\mu_{u}$, at which the ratios between the differences between the inverse fine structure constants for the three subgroups of the Standard Model group is the one predicted from our calculation of the quantum fluctuation corrections. In fact the three inverse fine structure constants shall lie on the number axis as the numbers $(2,13/5,3)$ corresponding to the subgroups $(SU(3),U(1),SU(2))$, where we have chosen the $SU(5)$-normalisation for the $U(1)$-finstructure constant. The relation is expressed in terms of the two independent differences, that can be formed. Let us , e.g., say
@@
$29.57\pm 0.06\%-\frac{5}{3}*59.00\pm 0.02\%+\frac{2}{3}*8.446\pm 0.6\%$ $=$ $\frac{-19/6-5/3*41/10+2/3*(-7)}{2\pi}*$
* $\ln\frac{\mu_{u}}{M_{Z}}$
@@ 7 What Says Our Result About Original Variables?
$-\frac{b_{2}-b_{3}}{b_{2}-5/3*b_{1}+2/3b_{3}}*$
$*(1/\alpha_{2}(M_{Z})-5/3*1/\alpha_{1\;SU(5)}(M_{Z})+2/3*1/\alpha_{3}(M_{Z})).$
@@ 8.2 Superfluous Case Difference 1 to 3
Although it is just related to the two foregoing let us explicitely for check also calculate what the our requirement for the difference 1to 3 means
@@ 9.1 More Thoughts on the Critical Coupling, Unified Coupling
Thinking a bit deeper: We should really not take a formula for the $SU(5)$ critical coupling without correction, because we have been claiming all through the article that in our model the $SU(5)$ symmetry and all its degrees of freedom do not exist. Rather we should look for correcting the number for the critical $\alpha_{5}$ to the critial coupling for the lattice standard Model group coupling:
@@ 10 Crude Second Order Calculation
One unneccesary ignorance of second order terms, which are expected to be of the order $(1/10)^{2}$ times the main term is, that we above let the $\alpha$ appearing as a factor in the $<H^{2}>$’s cancel with the $1/\alpha$ whichever among the $1/\alpha_{i}$’s we meet. Actually it was tempting to think, that by using this lucky trick of getting rid of the parameters in the estimate of our corrections, we were likely actually to get a better result with respect to agreeing for our calculations. Once we look for accuracies of the second order, there may be more corrections, such as the $H$ distribtuion being not even just Gaussian, and the whole program of doing the second order deserves a further article. Here we shall only make a very crude attempt to estimate the effect of seeing what $\alpha$ (among the three one say) comes into which of the fluctuations $<H^{2}>$. We shall make the assumption that the $\alpha$ to be used for an $H_{i}$ where it is the fuctuation in one of the basis vectors for the subgroup $i$ of the $SU(5)$, is $\alpha_{i}$. Then we see from table 1 that we have had relativley good luck by letting the two $\alpha$’s cancel each other, because the mostly contributing $<H_{i}^{2}>$ to the correction for the inverse fine structure constant $1/\alpha_{j}$, for the standard model subgroup denoted $j$ is actually mostly $i=j$ itself. In fact e.g. according to the table 1 the correction to
@@
For the $U(1)$ inverse fine structure constant the dominant contributuion to the corrections comes from the two nonabelian groups, i.e. from $H_{2}$ and $H_{3}$, but it has a bigger number from the $H_{1}$ than any of the other two groups, namely $7/30$. But since the $U(1)$ coupling correction is so mixed, to take all the same $\alpha$ is not so bad.
@@ 11 Speculative Relation to Planck Scale
A major problem and surprize comming, if one takes our suggestion of the truly existing lattice at the approximate or ours unification scale $\mu_{U}$ = $5.18*10^{13}GeV$ seriously is, that it suggests a “fundamental” scale quite different from the Planck scale. To seek a way out of this problem we propose to think of a fluctuating lattice even in size of the lattice constant in the sense that we speculate, that the general theory of relativity is still perturbatively treatable and rather well understood already - so that no completely speculated quatum gravity theory is needed at the $\mu_{U}$ scale - so that the whole lattice structure must be in a quantum superpostion state invariant under the reparametrization group from the general relativity. That is to say, with the philosophy, that there is very big quantum fluctuations in the gauge and taking the diffeomorphism of reparametrization symmetry as the gauge symmetry of general relativity, we must take it that the world is in a superposition of all the possible deformations of the lattice - needed for our model for the approximate GUT $SU(5)$ achieved by reparametrizations. That is to say, that in a typical component in this superpostion somewhere we find a very small lattice constant and somewherewe find a very big one, so that lattice cannot be exactly a Wilson one e.g.. But locally it could still be close to a Wilson lattice. Then of course the lattice constant value suggested by our parameter $\mu_{U}$ as lattice constant $a\approx 1/\mu_{U}$ could only be true in an average sense:
@@ 11.3 Ambiguity of Concept of Planck Energy Scale, Reduced ?
and has $\hbox{and has }\ln(\frac{M_{Z}}{m_{av.ch.fermions}})$ $=$ 6.3276.327 6.3276.327 (273)
@@
In fact we think, we can argue for, that this latter choice is not the correct one, because the $8\pi$ or $4\pi$ usually comes from the differnce in the coefficient to a Coulomb field and the charge appearing in the field theory action. When we have just used the fermion masses without any $4\pi$-like correction we associate it with the simple relation $m=g_{y}<\phi>$, while if I would like the Yukawa-field around the Higgs particle I would get a $1/(4\pi)$ factor in. So the simple masses correspond we could say to the Yukawa coupling $g_{y}$ being used for unit, and not the alternative $g_{y}/(4\pi)$. So to speak